# PerturBench: Benchmarking Machine Learning Models for Cellular Perturbation Analysis

### v3

@@ 4.3 Benchmarking Rules
Since model performance often varies with hyperparameters, we run hyperparameter optimization (HPO) for every model in each dataset, task and scenario to ensure accurate model comparisons. Specifically, we use optuna [3] with the default tree-structured Parzen estimator [7]. Each HPO run includes at least 60 trials with 6 trials running in parallel, and we select the best hyperparameters using a combination of the RMSE and RMSE rank metric, i.e. $\mathrm{RMSE}+0.1\cdot\mathrm{rank}_{\mathrm{RMSE}}$. We find this approach results in good overall performance. Additional details are in Appendix 10.6, along with the best hyperparameters for each model/dataset/task. For the best hyperparameter configuration, we run model training four additional times with different seeds to assess stability.
@@ 5 Experiments
In this section, we summarize the results of the covariate transfer and combo prediction tasks using different datasets, as well as the data scaling scenario. Additional results, including results for the Frangieh21 dataset, imbalanced data scenario, and implementation details can be found in Appendices 9 and 10.
Table 2: Results of the covariate transfer experiment measuring generalization across cell types in the Srivatsan20 dataset. Mean $\pm$ one standard deviation reported across 5 seeds. Best performance per metric is indicated in bold. The sme convention applies to subsequent tables.
Model Cosine RMSE Cosine RMSE
log fold change (LogFC) mean LogFC rank mean rank
CPA $0.38\pm 6\times 10^{-3}$ $0.020\pm 3\times 10^{-4}$ $0.15\pm 1\times 10^{-2}$ $0.16±\pm 7\times 10^{-3}$
CPA (noAdv) $0.40\pm 5\times 10^{-3}$ $0.020\pm 1\times 10^{-4}$ $\mathbf{0.09\pm 4\times 10^{-3}}$ $\mathbf{0.10\pm 5\times 10^{-3}}$
CPA (scGPT) $0.39\pm 9\times 10^{-3}$ $0.020\pm 3\times 10^{-4}$ $0.13\pm 2\times 10^{-2}$ $0.16\pm 2\times 10^{-2}$
SAMS-VAE $0.44\pm 1\times 10^{-3}$ $0.023\pm 8\times 10^{-5}$ $0.17\pm 1\times 10^{-2}$ $0.17\pm 1\times 10^{-2}$
SAMS-VAE (S) $\mathbf{0.53\pm 1\times 10^{-2}}$ $\mathbf{0.018\pm 3\times 10^{-4}}$ $0.12\pm 2\times 10^{-2}$ $0.13\pm 1\times 10^{-2}$
Biolord $0.18\pm 1\times 10^{-1}$ $0.086\pm 4\times 10^{-2}$ $0.37\pm 2\times 10^{-2}$ $0.35\pm 1\times 10^{-1}$
LA $0.45\pm 2\times 10^{-3}$ $\mathbf{0.018\pm 6\times 10^{-5}}$ $0.13\pm 4\times 10^{-3}$ $0.15\pm 3\times 10^{-3}$
LA (scGPT) $0.50\pm 4\times 10^{-3}$ $\mathbf{0.017\pm 1\times 10^{-4}}$ $0.13\pm 7\times 10^{-3}$ $0.14\pm 5\times 10^{-3}$
Decoder $0.35\pm 5\times 10^{-3}$ $\mathbf{0.018\pm 1\times 10^{-4}}$ $0.16\pm 1\times 10^{-2}$ $0.14\pm 7\times 10^{-3}$
Decoder (Cov) $0.30\pm 1\times 10^{-2}$ $0.023\pm 3\times 10^{-5}$ $0.47\pm 9\times 10^{-3}$ $0.50\pm 4\times 10^{-2}$
Linear $0.16\pm 1\times 10^{-2}$ $0.030\pm 5\times 10^{-4}$ $0.28\pm 5\times 10^{-3}$ $0.27\pm 2\times 10^{-3}$
5.1 Predicting Perturbation Effects Across Cell Types
We begin with the covariate transfer task and assess a model’s ability to predict the effects of drug treatment in cell types where the drugs have not been observed. For each cell line in the Srivatsan20 dataset, we held out 30% of the perturbations for validation and testing, ensuring that any held-out perturbations have been observed in the two other cell types.
The results are summarized in Table 2. Of the three published models, CPA performs best on the rank metrics, while SAMS-VAE is slightly worse. However, SAMS-VAE performs better on the fit based RMSE and cosine LogFC metrics. BioLord underperforms both models on all metrics.
@@ 5.1 Predicting Perturbation Effects Across Cell Types
Our decoder-only baseline which does not leverage perturbation information, Decoder (Cov), also performs well on cosine LogFC and RMSE fit, yet does no better than random on rank metrics. This suggests that Decoder (Cov) can find a single expression vector that achieves a decent fit to all perturbations in a given cell type, highlighting the need for our rank metrics that assess whether models can correctly order perturbations. Appendix 9.3 contains a more detailed assessment, where we are able to establish that SAMS-VAE (S) as well as our baselines, are less prone to mode collapse, based on analysis in the Srivatsan20 and Frangieh21 datasets.
@@
We also benchmark model performance with the Frangieh21 dataset, where we assess models’ ability to transfer perturbation effects from simpler cell systems to a more complex co-culture system. We have found similar trends in model performance, with more details in Appendix 9.1
Our simpler baselines, in particular the latent additive (LA) model, offer competitive performance at a fraction of training cost. These baseline models also scale well with larger datasets that contain more complex covariates, as shown in the Section on 5.3 with the McFalineFigueroa23 and Jiang24 datasets, where they outperform the more complex VAEs. Thus, it is important to evaluate models across a diverse set of datasets, which we offer in our benchmark.
@@ 5.2 Predicting Combinatorial Gene Over-expression Effects
Next we discuss model performance on the combo prediction task using the Norman19 dataset, which contains both single and dual genetic perturbations. The models see all single perturbations and 30% of the dual perturbations during training, with the remaining 70% of duals held out for validation and testing. We summarize the results in Table 3.
@@
Model Cosine RMSE Cosine RMSE
LogFC mean LogFC rank mean rank
CPA $0.76\pm 4\times 10^{-3}$ $0.046\pm 5\times 10^{-4}$ $0.0072\pm 2\times 10^{-3}$ $0.019\pm 3\times 10^{-3}$
CPA (noAdv) $0.77\pm 1\times 10^{-2}$ $0.046\pm 1\times 10^{-3}$ $\mathbf{0.0057\pm 3\times 10^{-3}}$ $0.017\pm 3\times 10^{-3}$
CPA (scGPT) $0.70\pm 2\times 10^{-2}$ $0.053\pm 1\times 10^{-3}$ $0.025\pm 6\times 10^{-3}$ $0.036\pm 3\times 10^{-3}$
SAMS-VAE $0.45\pm 2\times 10^{-2}$ $0.084\pm 7\times 10^{-4}$ $0.021\pm 5\times 10^{-3}$ $0.026\pm 2\times 10^{-3}$
SAMS-VAE (S) $\mathbf{0.78\pm 6\times 10^{-3}}$ $0.047\pm 2\times 10^{-3}$ $0.019\pm 5\times 10^{-3}$ $0.030±\pm 7\times 10^{-3}$
Biolord $0.41\pm 2\times 10^{-2}$ $0.086\pm 6\times 10^{-4}$ $0.027\pm 1\times 10^{-3}$ $0.028\pm 1\times 10^{-3}$
GEARS $0.44\pm 5\times 10^{-3}$ $0.069\pm 1\times 10^{-3}$ $0.051\pm 1\times 10^{-2}$ $0.055\pm 6\times 10^{-3}$
LA $\mathbf{0.79\pm 1\times 10^{-2}}$ $\mathbf{0.043\pm 4\times 10^{-4}}$ $\mathbf{0.005\pm 2\times 10^{-3}}$ $\mathbf{0.014\pm 1\times 10^{-3}}$
LA (scGPT) $0.77\pm 4\times 10^{-3}$ $\mathbf{0.044\pm 4\times 10^{-4}}$ $0.0085\pm 1\times 10^{-3}$ $\mathbf{0.013\pm 2\times 10^{-3}}$
Decoder $0.73\pm 2\times 10^{-2}$ $\mathbf{0.043\pm 3\times 10^{-4}}$ $0.017\pm 6\times 10^{-3}$ $\mathbf{0.014\pm 4\times 10^{-4}}$
Linear $0.60\pm 2\times 10^{-2}$ $0.057\pm 3\times 10^{-3}$ $0.035\pm 4\times 10^{-3}$ $0.016\pm 8\times 10^{-4}$
@@ 5.3 Effect of Data Scaling
Table 4: Results of the Jiang24 experiment which contains 30 different cell states (6 cell lines & 5 cytokine treatments), with 70% of perturbations from 9 cell states held out for validation/testing. Results are reported as mean $\pm$ one standard deviation. Best performance per metric is indicated in bold.
Model Cosine RMSE Cosine RMSE
LogFC mean LogFC rank mean rank
CPA $0.60\pm 7\times 10^{-4}$ $\mathbf{0.015\pm 5\times 10^{-5}}$ $0.40\pm 9\times 10^{-3}$ $0.42\pm 1\times 10^{-2}$
CPA (noAdv) $0.60\pm 2\times 10^{-3}$ $\mathbf{0.015\pm 4\times 10^{-5}}$ $0.39\pm 1\times 10^{-2}$ $0.40\pm 1\times 10^{-2}$
SAMS-VAE $0.59\pm 3\times 10^{-3}$ $0.017\pm 2\times 10^{-4}$ $0.45\pm 9\times 10^{-3}$ $0.45\pm 6\times 10^{-3}$
SAMS-VAE (S) $0.57\pm 5\times 10^{-2}$ $0.016\pm 2\times 10^{-3}$ $0.42\pm 2\times 10^{-2}$ $0.42\pm 3\times 10^{-2}$
LA $0.47\pm 1\times 10^{-3}$ $\mathbf{0.015\pm 7\times 10^{-5}}$ $0.38\pm 6\times 10^{-3}$ $0.38\pm 7\times 10^{-3}$
Decoder $\mathbf{0.64\pm 8\times 10^{-4}}$ $\mathbf{0.015\pm 3\times 10^{-5}}$ $\mathbf{0.32\pm 8\times 10^{-3}}$ $\mathbf{0.32\pm 5\times 10^{-3}}$
Linear $0.17\pm 8\times 10^{-5}$ $0.038\pm 5\times 10^{-5}$ $0.34\pm 2\times 10^{-3}$ $0.43\pm 1\times 10^{-3}$
In this section, we report the data scaling results, assessing whether models can take advantage of additional training data to better generalize perturbation effects across biological states. We use the McFalineFigueroa23 dataset that contains 3 cell lines with 5 chemical perturbations and 525 genetic perturbations. We treat unique cell line and chemical perturbation as a separate biological state, resulting in 15 total states. To test whether adding biological states improves performance, we construct nested subsets of the dataset ($small\subset medium\subset full$), all sharing the same validation and test sets. Each of the subsets contains more biological states (details in Appendix 10.4).
@@
The results are summarized in Table 4 where the decoder-only model performs the best for all metrics considered. Here, none of the published models perform on-par with our simple baselines, particularly on the rank metrics which indicates these more sophisticated models are unable to generate perturbation-specific responses. Again, this aligns with our earlier observation that simpler architectures with fewer modeling assumptions benefit more from additional training data.
@@ 6 Discussion
Our study shows that for predicting average perturbation effects, there is no model, or even class of models, that is clearly better than others. In smaller datasets, methods such as CPA (noAdv) and SAMS-VAE (S) exhibit the best performance. However, these methods do not scale well in larger datasets, where they are outperformed by our latent additive or decoder only models using the matching approach. The variance in performance across datasets highlights the need to develop more versatile and universally robust models and we hope that our perturbation framework will be helpful for developing and benchmarking these future models.
@@
We also find that ablating the adversarial component in CPA and the sparsity-inducing component in SAMS-VAE improve performance over the original models. Since it is highly probable that assumptions made and tested in one dataset may not hold true in another, comprehensive ablation experiments across diverse datasets and metrics as are provided in our framework will be a gateway to developing more versatile and robust model.
@@
Our simple decoder-only models with no gene expression input generally perform well and on Jiang24 dataset, significantly outperforms all other models. This potentially points to new architectures that emphasize the perturbation encoder, assuming the information present in the gene expression input may not be as rich as the perturbation and covariate information. But other possible explanations include inadequate control matching due to heterogeneity in the control cells not captured by covariate labels, as well as noise in either the single-cell RNA readouts or the perturbation labels. Specifically, CRISPR knockdown or knockout can be inefficient, resulting in some cells receiving the correct guide RNA but not having a robust gene knockdown or knockout [33].
@@
The performance on the larger datasets and tasks, that is McFalineFigueroa23 (data scaling) and Jiang24, are unsatisfactory compared to what was observed on the smaller datasets. As an example, on the data scaling task, none of the models got rank metric below $0.4$ (where $0.5$ would be achieved by a random predictor). This could be due to both the intrinsic difficulty of the task, and the amount of measurement or biological noise present in a dataset. For the data scaling task, most models performed better with more data, potentially suggesting perturbation models follow scaling laws [26].
Limitations
@@
Given the diversity of implementations among public models, we aimed to assess the core components of each model. Thus, our benchmarking results should be interpreted as an assessment of how these core components perform rather than a perfectly accurate recreation of the public model implementation (see Appendix 10.5 for details).
Since our benchmarking metrics are defined on a population level, they may not fully capture heterogeneity in the perturbation response among cells. We’ve implemented distributional metrics, such as maximum mean discrepancy (MMD) with RBF kernel and energy distance, as well as common differential expression metrics, such as log p-values, to better capture response heterogeneity and hope to use these metrics in future studies [20, 53].
@@
We provide three main components: datasets and dataloaders, a model development framework, and an evaluation API with metrics (Appendix 7). Each component can be used together or individually. For example, a user who just wants to benchmark predictions generated by an existing model can use the evaluation API. Whereas a user who wants a develop a new model with our model framework can use the entire codebase. Each component is extensible, making it easy for users to add new datasets, models, and metrics. The codebase can be found on GitHub at https://github.com/altoslabs/perturbench/.

### v4

@@ 4.3 Benchmarking Rules
Since model performance often varies with hyperparameters, we run hyperparameter optimization (HPO) for every model in each dataset, task and scenario to ensure accurate model comparisons. Specifically, we use optuna [4] with the default tree-structured Parzen estimator [8]. Each HPO run includes at least 60 trials with 6 trials running in parallel, and we select the best hyperparameters using a combination of the RMSE and RMSE rank metric, i.e. $\mathrm{RMSE}+0.1\cdot\mathrm{rank}_{\mathrm{RMSE}}$. We find this approach results in good overall performance. Additional details are in Appendix D.6, along with the best hyperparameters for each model/dataset/task. For the best hyperparameter configuration, we run model training four additional times with different seeds to assess stability.
@@ 5 Experiments
In this section, we summarize the results of the covariate transfer and combo prediction tasks using different datasets, as well as the data scaling scenario. Additional results, including results for the Frangieh21 and OP3 dataset, imbalanced data scenario, and implementation details can be found in Appendices CD, and D.7.
Table 2: Results of the covariate transfer experiment measuring generalization across cell lines in the Srivatsan20 dataset. Mean $\pm$ one standard deviation reported across 5 seeds. Best performance per metric is indicated in bold. The same convention applies to subsequent tables.
Model Cosine LogFC Cosine LogFC rank MMD PCA DEG recall
CPA $0.38\pm 6\times 10^{-3}$ $0.15\pm 1\times 10^{-2}$ $53\pm 4\times 10^{-3} 0.0$$007\pm 2\times 10^{-3}$
[…]
CPA (scGPT) $0.39\pm 9\times 10^{-3}$ $0.13\pm 2\times 10^{-2}$ - -
SAMS-VAE $0.44\pm 1\times 10^{-3}$ $0.17\pm 1\times 10^{-2}$ $69\pm 1\times 10^{-2} 0.0$$000\pm 1\times 10^{-4}$
SAMS-VAE (S) $\mathbf{0.53\pm 1\times 10^{-2}}$ $0.12\pm 2\times 10^{-2}$ $79\pm 1\times 10^{-2} 0.0$$000\pm 0$
Biolord $0.18\pm 1\times 10^{-1}$ $0.37\pm 2\times 10^{-2}$ $4.3\pm 4\times 10^{0}$ $0.000\pm 1\times 10^{-4}$
LA $0.45\pm 2\times 10^{-3}$ $0.13\pm 4\times 10^{-3}$ $2.0\pm 2\times 10^{-1}$ $0.000\pm 0$
LA (scGPT) $0.50\pm 4\times 10^{-3}$ $0.13\pm 7\times 10^{-3}$ - -
Decoder $0.35\pm 5\times 10^{-3}$ $0.16\pm 1\times 10^{-2}$ $1.9\pm 5\times 10^{-3}$ -
Decoder (Cov) $0.30\pm 1\times 10^{-2}$ $0.47\pm 9\times 10^{-3}$ - -
Linear $0.16\pm 1\times 10^{-2}$ $0.28\pm 5\times 10^{-3}$ $76\pm 9\times 10^{-4} 0.0$$004\pm 3\times 10^{-4}$
5.1 Predicting Perturbation Effects Across Cell Lines
We begin with the covariate transfer task and assess each model’s ability to predict the effects of drug treatment in cell lines where the drugs have not been observed. For each cell line in the Srivatsan20 dataset, we held out 30% of the perturbations for validation and testing, ensuring that any held-out perturbations have been observed in the two other cell lines. The results are summarized in Table 2. Of the three published models, CPA performs best on the rank metrics, while SAMS-VAE is slightly worse. However, SAMS-VAE performs better on the fit based RMSE and cosine LogFC metrics. BioLord underperforms both models on all metrics.
@@ 5.1 Predicting Perturbation Effects Across Cell Lines
Our Decoder-Only baseline which does not leverage perturbation information, Decoder (Cov), also performs well on cosine LogFC and RMSE fit, yet does no better than random on rank metrics. This suggests that Decoder (Cov) can find a single expression vector that achieves a decent fit to all perturbations in a given cell line, highlighting the need for our rank metrics that assess whether models can correctly order perturbations. Appendix C.7 contains a more detailed assessment, where we are able to establish that SAMS-VAE (S) as well as our baselines, are less prone to mode collapse, based on analysis in the Srivatsan20 and Frangieh21 datasets.
@@
We see large differences in model performance on the MMD metric computed in the PCA space (MMD PCA). Models using autoencoder architectures (CPA and SAMS-VAE), tend to have better performance than our baseline models (Latent Additive, Decoder Only). All models performed poorly on our differentially expressed gene recall (DEG recall) metric. Additional metrics for Srivatsan20 can be found in Table 5.
Since the Srivatsan20 dataset contained mostly cancer cell lines, we also benchmarked models on the Frangieh21 and OP3 datasets, which contain melanoma cells in co-culture with immune cells and primary peripheral blood monocyte cells (PBMCs) respectively. With the Frangieh21 dataset, we assess models’ ability to transfer genetic perturbation effects from simpler cell systems to a more complex co-culture system, finding similar trends in model performance, with more details in Appendix C.3.
With the OP3 dataset, we assess models’ ability to transfer small molecule effects across heterogeneous PBMC cell types, again finding similar trends in model performance. The gap between CPA/SAMS-VAE and the Latent Additive/Decoder models’ performance on the MMD PCA metric was larger, suggesting that the OP3 dataset contains greater within-cell-type heterogeneity compared to the Srivatsan20 or Frangieh21 datasets. Additional details can be found in Appendix C.4
Our simpler baselines, in particular the Latent Additive (LA) model, offer competitive performance at a fraction of training cost. These baseline models also scale well with larger datasets that contain more complex covariates, as shown in the Section 5.3 with the McFalineFigueroa23 and Jiang24 datasets, where they outperform the more complex VAEs. However, they underperform more complex models when evaluated on the MMD metrics, suggesting they only learn an average response to perturbation. Thus, it is important to evaluate models across a diverse set of datasets and metrics, which we offer in PerturBench.
@@ 5.2 Predicting Combinatorial Gene Over-expression Effects
Next we discuss model performance on the combo prediction task using the Norman19 dataset, which contains both single and dual genetic perturbations. The models train on all single perturbations and 30% of the dual perturbations, with the remaining 70% of duals held out for validation and testing. We summarize the results in Table 3.
@@
The autoencoder architectures again outperform the baseline models in the MMD PCA metric. However, the gap is smaller than the gap observed in the Srivatsan20, Frangieh21, or OP3 datsets, suggesting less within-cell-line heterogeneity in the Norman19 dataset.
@@
Model Cosine logFC Cosine LogFC rank MMD PCA DEG recall
CPA $76\pm 4\times 10^{-3} 0.0$$0072\pm 2\times 10^{-3}$ $2.2\pm 2\times 10^{-2}$ $0.032\pm 4\times 10^{-3}$
[…]
CPA (scGPT) $70\pm 2\times 10^{-2} 0.0$$025\pm 6\times 10^{-3} - -$
SAMS-VAE $45\pm 2\times 10^{-2} 0.0$$021\pm 5\times 10^{-3}$ $1.9\pm 3\times 10^{-2}$ $0.000\pm 0$
SAMS-VAE (S) $\mathbf{0.78\pm 6\times 10^{-3}}$ $0.019\pm 5\times 10^{-3}$ $\mathbf{0.74\pm 5\times 10^{-2}}$ $0.028\pm 6\times 10^{-3}$
Biolord $41\pm 2\times 10^{-2} 0.0$$027\pm 1\times 10^{-3}$ $1.6\pm 5\times 10^{-3}$ $0.000\pm 0$
GEARS $44\pm 5\times 10^{-3} 0.0$$051\pm 1\times 10^{-2} - -$
LA $\mathbf{0.79\pm 1\times 10^{-2}}$ $\mathbf{0.005\pm 2\times 10^{-3}}$ $3.2\pm 6\times 10^{-3}$ $0.000\pm 3\times 10^{-4}$
LA (scGPT) $77\pm 4\times 10^{-3} 0.0$$0085\pm 1\times 10^{-3} - -$
Decoder $73\pm 2\times 10^{-2} 0.0$$017\pm 6\times 10^{-3}$ $3.2\pm 4\times 10^{-3}$ -
Linear $60\pm 2\times 10^{-2} 0.0$$035\pm 4\times 10^{-3}$ $1.2\pm 4\times 10^{-2}$ $0.018\pm 2\times 10^{-3}$
@@ 5.3 Effect of Data Scaling
Table 4: Results of the Jiang24 experiment which contains 30 different biological states (6 cell lines & 5 cytokine treatments), with 70% of perturbations from 9 states held out for validation/testing. Results are reported as mean $\pm$ one standard deviation. Best performance per metric is indicated in bold.
Model Cosine logFC Cosine LogFC rank MMD PCA DEG recall
CPA $0.60\pm 7\times 10^{-4}$ $0.40\pm 9\times 10^{-3}$ $2.5\pm 4\times 10^{-3}$ $0.004\pm 2\times 10^{-3}$
CPA (noAdv) $0.60\pm 2\times 10^{-3}$ $0.39\pm 1\times 10^{-2}$ $2.5\pm 3\times 10^{-3}$ $0.005\pm 6\times 10^{-4}$
SAMS-VAE $0.59\pm 3\times 10^{-3}$ $0.45\pm 9\times 10^{-3}$ $\mathbf{0.32\pm 5\times 10^{-3}}$ $0.000\pm 4\times 10^{-5}$
SAMS-VAE (S) $0.57\pm 5\times 10^{-2}$ $0.42\pm 2\times 10^{-2}$ $1.6\pm 1\times 10^{0}$ $0.002\pm 1\times 10^{-4}$
LA $0.47\pm 1\times 10^{-3}$ $0.38\pm 6\times 10^{-3}$ $2.6\pm 2\times 10^{-3}$ $0.001\pm 5\times 10^{-4}$
Decoder $\mathbf{0.64\pm 8\times 10^{-4}}$ $\mathbf{0.32\pm 8\times 10^{-3}}$ $2.6\pm 2\times 10^{-3}$ -
Linear $0.17\pm 8\times 10^{-5}$ $0.34\pm 2\times 10^{-3}$ $1.3\pm 2\times 10^{-3}$ $0.003\pm 2\times 10^{-4}$
In this section, we report the data scaling results, assessing whether models can take advantage of additional training data to better generalize perturbation effects across biological states. We use the McFalineFigueroa23 dataset that contains 3 cell lines with 5 chemical perturbations and 525 genetic perturbations. We treat unique cell line and chemical perturbation as a separate biological state, resulting in 15 total states. To test whether adding biological states improves performance, we construct nested subsets of the dataset ($small\subset medium\subset full$), all sharing the same validation and test sets. Each of the subsets contains more biological states (details in Appendix D.4).
@@
The results are summarized in Table 4 where the Decoder-Only model performs the best on average response and SAMS-VAE performs best on the MMD PCA metric. Here, none of the published models perform on-par with our simple baselines, particularly on the rank metrics which indicates these more sophisticated models are unable to generate perturbation-specific responses. There is less of a gap between the more complex autoencoder architectures and the baseline models in the MMD PCA metric, suggesting less within-covariate heterogeneity compared to other datasets. These results align with our earlier observation that simpler architectures with fewer modeling assumptions benefit more from additional training data.
@@ 6 Discussion
Our study shows that for predicting average perturbation effects, there is no model, or even class of models, that is clearly better than others. In smaller datasets, methods such as CPA (noAdv) and SAMS-VAE (S) exhibit the best performance. However, these methods do not scale well in larger datasets, where they are outperformed by our Latent Additive or Decoder-Only models using the matching approach. The variance in performance across datasets highlights the need to develop more versatile and universally robust models and we hope that our perturbation framework will be helpful for developing and benchmarking these future models.
In terms of predicting the distribution of perturbation responses, as measured by the MMD metrics, autoencoder models such as CPAand SAMS-VAE outperform the Latent Additive and Decoder-Only baselines. The gap in MMD performance depends on the dataset, suggesting that some datasets have less heterogeneity within biological states. More recent model architectures such as STATE and CellFlow may be able to predict the full distribution of perturbation responses while also maintaining strong performance on average effect metrics [2, 28]
@@
We also find that ablating the adversarial component in CPA and the sparsity-inducing component in SAMS-VAE improve performance over the original models. Since it is highly probable that assumptions made and tested in one dataset may not hold true in another, comprehensive ablation experiments across diverse datasets and metrics as are provided in Perturbench will be a essential to developing more versatile and robust model architectures.
@@
Our simple Decoder-Only models with no gene expression input generally perform well, outperforming all other models on the Jiang24 dataset. This potentially points to new architectures that emphasize the perturbation encoder, assuming the information present in the gene expression input may not be as rich as the perturbation and covariate information. Other possible explanations include inadequate control matching due to heterogeneity in the control cells not captured by covariate labels, as well as noise in the single-cell RNA readouts or the perturbation labels. Specifically, some cells with a perturbation label may not have a robust target gene knockdown in CRISPR experiments [32].
@@
Models perform worse on the larger datasets and tasks, specifically McFalineFigueroa23 (data scaling) and Jiang24. For example, on the data scaling task, no model achieved a rank metric below $0.4$ (where $0.5$ is random predictor). This could be due to both the intrinsic difficulty of the task, and the amount of measurement or biological noise present in a dataset. For the data scaling task, most models performed better with more data, potentially suggesting perturbation models follow scaling laws [25].
@@
Given the diversity of implementations among public models, we aimed to assess the core components of each model. Thus, our benchmarking results should be interpreted as an assessment of how these core components perform rather than a perfectly accurate recreation of the public model implementation (see Appendix D.5 for details).
@@
We provide three main components: datasets and dataloaders, a model development framework, and an evaluation API with metrics (Appendix A). Each component can be used together or individually. For example, a user who just wants to benchmark predictions generated by an existing model can use the evaluation API. Whereas a user who wants to develop a new model with our model framework can use the entire codebase. Each component is extensible, making it easy for users to add new datasets, models, and metrics. The codebase can be found on GitHub at https://github.com/altoslabs/perturbench/.